# labelinglabel

### v1

@@ 6.1 A Phantom Nemesis?
It is presently unknown whether the observed accelerated expansion of our Universe can be generated by some fundamental phantom fields. Even a question of whether such fields might exist at all is a subject of rather heated debates. What is relevant for us here is one undeniable fact: the Big Rip singularities, deeply associated with such phantom fields, are utterly unreconcilable with the very method of quantization via MIU. To understand why, let us first recall that the thermodynamics of the phantoms require them to possess negative absolute temperatures [40], see also [41]. Consider, for example, a universe with scale factor $a$, filled by a minimally coupled scalar field $\phi$. A density of the kinetic energy is proportional to a density of temperature $T$, so
@@
where $K>0$ is a constant of integration. A phantom field has $w<-1$, so the temperature (62) ought to be negative. On the other hand, an entropy of a phantom field can be shown to remain positive [40]  [^13] One can easily verify this by expressing $\rho=\rho_{0}a^{-3(w+1)}$ via the temperature (62) and then using the Second law of thermodynamics.. The most natural explanation of this discrepancy must lie in a fundamentally quantum mechanical nature of the phantoms. And, since arguably the strongest manifestation of a phantom field is a Big Rip singularity, it in turn must also be a purely quantum phenomenon. An alternative way to reach this conclusion lies in observing the growth of an energy density at a vicinity of Big Rip: once it reaches the Planck value $\rho_{{}_{PL}}=c^{5}\hbar^{-1}G^{-2}$, a correct description of the dynamics can no longer be classical and has to account for the quantum phenomena. Anyway, one way or another, we have apparently stumbled upon a serious problem. According to the MIU formalism, all quantum mechanical effects have but one source: a repelling quantum potential (14). And this potential will obviously vanish during the uninhibited growth of the scale factor as we approach the Big Rip singularity! Hence, according to the MIU model, the quantum mechanical effects will vanish just as we approach Big Rip. So, we end up with a nasty contradiction on our hands, which can only be resolved in two possible ways: (1) The MIU approach might be incorrect (needless to say, this would be a huge blow for the authors!) or, at the very least, it might require certain modifications to ensure the growth of the quantum effects near the Big Rip; (2) The MIU is correct, and the problem is alleviated because the quantum potential actively prevents the Big Rip singularities from arising just like it did with big bang/crunch singularities. Very fortunately for the authors, the very conditions we have used in Theorem 1 ensures that the latter is actually true!
Before we proceed to the crux of our proof, let us list everything we’ll need and will use. First of all, for simplicity we will stick to a unit system where $18\pi G/3=c=$ 1. Secondly, we will use the equation (16), (17) with the potential (14) and a boundary condition (15). Thirdly, the continuity condition (21) must also be taken into account. Finally, out of four conditions from Sec. 3 we will utilize the first three, (18), (19), (20), but not the fourth one – after all, one can hardly hope to vanquish the Big Rip singularities if they are not even given a chance to appear!
@@ 6.2 The Step-by-Step Proof of Theorem 2
[…]
@@
    and formally write the system $N-1$ equations (16) as:
@@
    Since $N-2$ functions (63) are known, one can derive $N-2$ new functions, henceforth known as ${\hat{a}}_{n}(t)$, such that the for all $n\geq 3$ the Wronskians $W(a_{n}(t),\,{\hat{a}}_{n}(t))=1$. Naturally, the functions will have the form
@@
[…]
@@
    What about ${\hat{a}}_{2}(t)$, which does not satisfy a Schrödinger equation? In order to find it we can substitute $N-2$ newly found functions ${\hat{a}}_{n}(t)$ plus the hitherto unknown ${\hat{a}}_{2}(t)$ into (14). The result will be the following nonlinear O.D.E. of second order:
@@
    By solving (66) for ${\hat{a}}_{2}(t)$ we would successfully define $N-1$ new functions ${\hat{a}}_{n}(t)$ for $n\geq 2$ – the last function ${\hat{a}}_{1}=a_{1}(t)$ we’ll assume to be the same zero universe (3)). Interestingly, the initial ordering of original scale factors $a_{n}(t)<a_{n+1}(t)$ $\forall\,n$ will generally be inverted for the “crowned” scale factors – i.e. we will usually have ${\hat{a}}_{n}(t)>{\hat{a}}_{n+1}(t)$ for $n>2$.
@@
    The $N-1$ functions ${\hat{a}}_{n}(t)$, $n\geq 2$, found on Step 3 will be the solutions of the same equations (16), (17) (as well as the system (64) with $a_{n}\to{\hat{a}}_{n}$ and the original “potentials” $v_{n}(t)$), but with the different fields of matter. In other words, they will correspond to $N-2$ new densities ${\hat{\rho}}_{n}(t)$ and new pressures ${\hat{p}}_{n}(t)$, whose forms can be derived as follows.
@@
      These ${\hat{p}}_{n}(t)$ we can then substitute into $N-2$ continuity equations – thus guaranteeing that (17) with newly “crowned” ${\hat{a}}_{n}$, ${\hat{\rho}}_{n}$ and ${\hat{p}}_{n}$ remains valid.
@@
      Integrating the resulting $N-2$ linear O.D.E.s of first order, we can find ${\hat{\rho}}_{n}(t)$. Substituting them back into the continuity equations (21), we also find ${\hat{p}}_{n}(t)$. Note, that it will suffice to solve just $N-3$ equations, since ${\hat{\rho}}_{3}(t)$ is then easily derived from (17).
@@ 7 Anisotropy, the Decoherence and the Initial Conditions Problem
Since this article is primarily concerned with the isotropic solutions, we will abstain from writing the complete Einstein equations for general anisotropic models. Instead we will provide one equation for a flat universe (for the sake of simplicity let us use the dimensional system where $18\pi G/3=c=$ 1), where $\rho$ is a zeroth component of the stress-energy tensor:
@@
This equation is well-suited for building up the generalized constraint conditions akin to (17). To do that one should first introduce $N$ universes with the metric (69) and coefficients $a_{n}(t)$, $b_{n}(t)$, $c_{n}(t)$, $n=1,...,N$ and then replace (70) with a sum over all those universes with an additional quantum potential $U(\{a_{1},b_{1},c_{1}\},..,\{a_{{}_{N}},b_{{}_{N}},c_{{}_{N}}\})$. The key question, of course, would the explicit form of this potential. For the homogenous and isotropic models we have used the potential in the form (14), originally derived in [7] for a one-dimensional problem. It seems logical to assume that the anisotropic potential can also be derived as a three-dimensional generalization of the potential from [7]. Naturally, it would be premature to draw any serious conclusion without knowing its explicit form, but we do know that it has to be positively defined. And it was this very property that led us to a conclusion about the absence of singularities in FLRW models. Hence, we have a serious possibility that the quantization by the MIU model equipped with the Barrow’s zero-universe might get rid of singularities even in anisotropic universes.
@@
A second major problem to ponder is the problem of proper initial conditions for the inflation. The theory of inflation is truly majestic for a number of reasons, not the least of whom is an readiness with which it explains the thermodynamical arrow of time. Unfortunately, the is a tiny cloud that brings an unwelcome shade upon the majesty – namely, the initial conditions required for the inflation must have radically low entropy. It is commonly believed that the solution for the problem lay dormant somewhere in the future theory of quantum gravity. Now, we would be first to fully admit that the MIU formalism is by no means that theory – it is at best a few introductory accords of the upcoming grand symphony of quantum gravity theory. Nevertheless, we can still try and use the MIU technique as a sort of a sonar to at least try and adumbrate the contours of the future theory. In particular, we can ask ourselves: does this approach bring anything new to the problem of the initial conditions? We’d argue that it does!
@@ A.2 MIW: Let there be interaction!
$\varphi\right>=\langle\psi|\varphi|\psi\rangle\approx\frac{1}{N}\sum\limits_{j=1}^{N}\varphi(\vec{x}_{j}),$ (85)
@@
where $\vec{p}_{j}$ is a momentum of $j$-th particle, $V(\vec{x}_{j})$ are classic potentials, and $U_{N}$ is a new potential of quantum interaction. The average energy $E\right>$ for our ensemble is simply
$E\right>=\frac{H_{N}}{N}.$ (88)
@@
[…]
@@
$E\right>\approx\frac{1}{N}\sum\limits_{j=1}^{N}\left(\frac{1}{2m}\left(\nabla S\right)^{2}+V+\frac{\hbar^{2}}{8m}\left(\frac{\nabla P}{P}\right)^{2}\right)\Bigg{|}_{\vec{q}=\vec{x}_{j}}.$
@@
$E\right>\approx\frac{1}{N}\sum\limits_{j=1}^{N}\left(\frac{(\vec{p}_{j})^{2}}{2m}+V(\vec{x}_{j})+\frac{\hbar^{2}}{8m}\left(\frac{\nabla P(x_{j})}{P(x_{j})}\right)^{2}\right).$ (91)
@@
[…]
@@ A.3 MIU: A World is Not Enough
Consider a homogenous isotropic universe with a given curvature $\kappa$ (equal to either 0 or ±$\pm 1$), filled by matter fields of density $\rho$ and pressure $p$. Such a universe is known to be governed by a single dynamic parameter – scale factor $a$, which satisfies the two Friedman-Lemaître-Robertson-Walker equations:
@@
$\mathcal{H}(a,{\mathbf{p}})=\frac{\mathbf{p}^{2}}{2}-\frac{1}{2}\rho(a)a^{2}=-\frac{\kappa}{2},$
@@
$\begin{split}\mathcal{H}_{N}=&\sum\limits_{n=1}^{N}\frac{\mathbf{p}_{n}^{2}}{2}-\frac{1}{2}\sum_{n=1}^{N}\rho_{n}(a_{n})a_{n}^{2}+\\ &+U(a_{1},..,a_{N})+\frac{1}{2}\sum_{n=1}^{N}\kappa_{n}=0,\end{split}$
@@
$\begin{split}L_{N}&=\frac{1}{2}\sum_{n=1}^{N}\dot{a}_{n}^{2}+\frac{1}{2}\sum_{n=1}^{N}\rho_{n}(a_{n})a_{n}^{2}-\\ &-\alpha^{2}\sum_{n=1}^{N}\left(\frac{1}{a_{n+1}-a_{n}}-\frac{1}{a_{n}-a_{n-1}}\right)^{2},\end{split}$
@@ Acknowledgments
The article was supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no. 075-02-2024-1430)

### v3

@@ 6.1 A Phantom Nemesis?
It is presently unknown whether the observed accelerated expansion of our Universe can be generated by some fundamental phantom fields. Even a question of whether such fields might exist at all is a subject of rather heated debates [40], [41], [42], [43], [44]. What is relevant for us here is one undeniable fact: the Big Rip singularities, deeply associated with such phantom fields, are utterly unreconcilable with the very method of quantization via MIU. To understand why, let us first recall that the thermodynamics of the phantoms require them to possess negative absolute temperatures [45], see also [46]. Consider, for example, a universe with scale factor $a$, filled by a minimally coupled scalar field $\phi$. A density of the kinetic energy is proportional to a density of temperature $T$, so
@@
where $K>0$ is a constant of integration. A phantom field has w<1w<-1, so the temperature (62) ought to be negative. On the other hand, an entropy of a phantom field can be shown to remain positive [45]  [^15] One can easily verify this by expressing $\rho=\rho_{0}a^{-3(w+1)}$ via the temperature (62) and then using the Second law of thermodynamics.. The most natural explanation of this discrepancy must lie in a fundamentally quantum mechanical nature of the phantoms. And, since arguably the strongest manifestation of a phantom field is a Big Rip singularity, it in turn must also be a purely quantum phenomenon. An alternative way to reach this conclusion lies in observing the growth of an energy density at a vicinity of Big Rip: once it reaches the Planck value $\rho_{{}_{PL}}=c^{5}\hbar^{-1}G^{-2}$, a correct description of the dynamics can no longer be classical and has to account for the quantum phenomena. Anyway, one way or another, we have apparently stumbled upon a serious problem. According to the MIU formalism, all quantum mechanical effects have but one source: a repelling quantum potential (14). And this potential will obviously vanish during the uninhibited growth of the scale factor as we approach the Big Rip singularity! Hence, according to the MIU model, the quantum mechanical effects will vanish just as we approach Big Rip. So, we end up with a nasty contradiction on our hands, which can only be resolved in two possible ways: (1) The MIU approach might be incorrect (needless to say, this would be a huge blow for the authors!) or, at the very least, it might require certain modifications to ensure the growth of the quantum effects near the Big Rip; (2) The MIU is correct, and the problem is alleviated because the quantum potential actively prevents the Big Rip singularities from arising just like it did with big bang/crunch singularities. Very fortunately for the authors, the very conditions we have used in Theorem 1 ensures that the latter is actually true!
Before we proceed to the crux of our proof, let us list everything we’ll need and will use. First of all, for simplicity we will stick to a unit system where $8\pi G/3=c=1$. Secondly, we will use the equation (16), (17) with the potential (14) and a boundary condition (15). Thirdly, the continuity condition (21) must also be taken into account. Finally, out of four conditions from Sec. 3 we will utilize the first three, (18), (19), (20), but not the fourth one – after all, one can hardly hope to vanquish the Big Rip singularities if they are not even given a chance to appear!
@@ 6.2 The Step-by-Step Proof of Theorem 2
    Let’s additionally assume that we have managed to find the exact solutions of (16), (17), i.e. that there exist (N1)(N-1) time-dependent functions $a_{n}(t)$ (n=2, 3,,Nn=2,\,3,...,\,N) and 2(N2)2(N-2) time-dependent functions $\rho_{n}(t)$, $p_{n}(t)$, with n=3,,Nn=3,...,\,N (recall that the universe with $n=1$ is a Barrow’s zero-universe (3), whereas the universe with $n=2$ is purely quantum and has no fields of matter). This means that we can define N2N-2 ($n\geq 3$) time-dependent functions $v_{n}(t)$:
@@
    and formally write the system N1N-1 equations (16) as:
@@
    Since N2N-2 functions (63) are known, one can derive N2N-2 new functions, henceforth known as ${\hat{a}}_{n}(t)$, such that the for all $n\geq 3$ the Wronskians $W(a_{n}(t),\,{\hat{a}}_{n}(t))=1$. Naturally, the functions will have the form
@@
    and they will also be solutions of Schödinger equations (64) with the same “potentials” $v_{n}(t)$, but linearly independent of $a_{n}(t)$ [^17] Generally speaking, if $y_{1}(x)$ is a non-trivial partial solution of ordinary linear O.D.E. $y^{\prime\prime}+p(x)y^{\prime}+q(x)y=0$ where $p(x)$ is a continuous function, then the linearly independent solution $y_{2}(x)$ of the same equation will have a form: $y_{2}=y_{1}(x)\int{\frac{\mathrm{e}^{-\int{p(x)dx}}}{y_{1}^{2}(x)}dx}.$ It is easy to check that in this case the Wronskian $W(y_{1},y_{2})=1$.
@@
    What about ${\hat{a}}_{2}(t)$, which does not satisfy a Schrödinger equation? In order to find it we can substitute N2N-2 newly found functions ${\hat{a}}_{n}(t)$ plus the hitherto unknown ${\hat{a}}_{2}(t)$ into (14). The result will be the following nonlinear O.D.E. of second order:
@@
    By solving (66) for ${\hat{a}}_{2}(t)$ we would successfully define N1N-1 new functions ${\hat{a}}_{n}(t)$ for $n\geq 2$ – the last function ${\hat{a}}_{1}=a_{1}(t)$ we’ll assume to be the same zero universe (3)). Interestingly, the initial ordering of original scale factors $a_{n}(t)<a_{n+1}(t)$ $\forall\,n$ will generally be inverted for the “crowned” scale factors – i.e. we will usually have ${\hat{a}}_{n}(t)>{\hat{a}}_{n+1}(t)$ for $n>2$.
@@
    The N1N-1 functions ${\hat{a}}_{n}(t)$, $n\geq 2$, found on Step 3 will be the solutions of the same equations (16), (17) (as well as the system (64) with $a_{n}\to{\hat{a}}_{n}$ and the original “potentials” $v_{n}(t)$), but with the different fields of matter. In other words, they will correspond to N2N-2 new densities ${\hat{\rho}}_{n}(t)$ and new pressures ${\hat{p}}_{n}(t)$, whose forms can be derived as follows.
@@
      These ${\hat{p}}_{n}(t)$ we can then substitute into N2N-2 continuity equations – thus guaranteeing that (17) with newly “crowned” ${\hat{a}}_{n}$, ${\hat{\rho}}_{n}$ and ${\hat{p}}_{n}$ remains valid.
@@
      Integrating the resulting N2N-2 linear O.D.E.s of first order, we can find ${\hat{\rho}}_{n}(t)$. Substituting them back into the continuity equations (21), we also find ${\hat{p}}_{n}(t)$. Note, that it will suffice to solve just N3N-3 equations, since ${\hat{\rho}}_{3}(t)$ is then easily derived from (17).
@@ 7 Anisotropy, the Decoherence and the Initial Conditions Problem
Since this article is primarily concerned with the isotropic solutions, we will abstain from writing the complete Einstein equations for general anisotropic models. Instead we will provide one equation for a flat universe (for the sake of simplicity let us use the dimensional system where $8\pi G/3=c=1$), where $\rho$ is a zeroth component of the stress-energy tensor:
@@
This equation is well-suited for building up the generalized constraint conditions akin to (17). To do that one should first introduce $N$ universes with the metric (69) and coefficients $a_{n}(t)$, $b_{n}(t)$, $c_{n}(t)$, n=1,,Nn=1,...,N and then replace (70) with a sum over all those universes with an additional quantum potential $U(\{a_{1},b_{1},c_{1}\},..,\{a_{{}_{N}},b_{{}_{N}},c_{{}_{N}}\})$. The key question, of course, would the explicit form of this potential. For the homogenous and isotropic models we have used the potential in the form (14), originally derived in [1] for a one-dimensional problem. It seems logical to assume that the anisotropic potential can also be derived as a three-dimensional generalization of the potential from [1]. Naturally, it would be premature to draw any serious conclusion without knowing its explicit form, but we do know that it has to be positively defined. And it was this very property that led us to a conclusion about the absence of singularities in FLRW models. Hence, we have a serious possibility that the quantization by the MIU model equipped with the Barrow’s zero-universe might get rid of singularities even in anisotropic universes.
@@
A second major problem to ponder is the problem of proper initial conditions for the inflation. The theory of inflation is truly majestic for a number of reasons, not the least of whom is an readiness with which it explains the thermodynamical arrow of time. Unfortunately, the is a tiny cloud that brings an unwelcome shade upon the majesty – namely, the initial conditions required for the inflation must have radically low entropy [5], [6]. It is commonly believed that the solution for the problem lay dormant somewhere in the future theory of quantum gravity. Now, we would be first to fully admit that the MIU formalism is by no means that theory – it is at best a few introductory accords of the upcoming grand symphony of quantum gravity theory. Nevertheless, we can still try and use the MIU technique as a sort of a sonar to at least try and adumbrate the contours of the future theory. In particular, we can ask ourselves: does this approach bring anything new to the problem of the initial conditions? We’d argue that it does!
@@ A.2 MIW: Let there be interaction!
$\left<\varphi\right>=\langle\psi|\varphi|\psi\rangle\approx\frac{1}{N}\sum\limits_{j=1}^{N}\varphi(\vec{x}_{j}),$ (85)
@@
where $\vec{p}_{j}$ is a momentum of $j$-th particle, $V(\vec{x}_{j})$ are classic potentials, and $U_{N}$ is a new potential of quantum interaction. The average energy $\left<E\right$> for our ensemble is simply
$\left<E\right>=\frac{H_{N}}{N}.$ (88)
@@
$\left<E\right>=\int dq~|\psi|^{2}\left(\frac{1}{2m}\left(\nabla S\right)^{2}+V+\frac{\hbar^{2}}{8m}\left(\frac{\nabla P}{P}\right)^{2}\right),$ (89)
@@
$\left<E\right>\approx\frac{1}{N}\sum\limits_{j=1}^{N}\left(\frac{1}{2m}\left(\nabla S\right)^{2}+V+\frac{\hbar^{2}}{8m}\left(\frac{\nabla P}{P}\right)^{2}\right)\Bigg{|}_{\vec{q}=\vec{x}_{j}}.$
@@
$\left<E\right>\approx\frac{1}{N}\sum\limits_{j=1}^{N}\left(\frac{(\vec{p}_{j})^{2}}{2m}+V(\vec{x}_{j})+\frac{\hbar^{2}}{8m}\left(\frac{\nabla P(x_{j})}{P(x_{j})}\right)^{2}\right).$ (91)
@@
$\begin{split}\frac{1}{N}\sum\limits_{j=1}^{N}\phi(x_{j})&\approx\sum\limits_{j=1}^{N}\int\limits_{x_{j}}^{x_{j+1}}dqP(x_{j})\phi(x_{j})\\ &=\sum\limits_{j=1}^{N}(x_{j+1}-x_{j})P(x_{j})\phi(x_{j}).\end{split}$ (94)
@@ A.3 MIU: A World is Not Enough
Consider a homogenous isotropic universe with a given curvature $\kappa$ (equal to either 0 or $\pm 1$), filled by matter fields of density $\rho$ and pressure $p$. Such a universe is known to be governed by a single dynamic parameter – scale factor $a$, which satisfies the two Friedman-Lemaître-Robertson-Walker equations:
@@
$\mathcal{H}(a,{\mathbf{p}})=\frac{\mathbf{p}^{2}}{2}-\frac{4\pi G}{3}\rho(a)a^{2}=-\frac{\kappa c^{2}}{2},$
@@
$\begin{split}\mathcal{H}_{N}=&\sum\limits_{n=1}^{N}\frac{\mathbf{p}_{n}^{2}}{2}-\frac{4\pi G}{3}\sum_{n=1}^{N}\rho_{n}(a_{n})a_{n}^{2}+\\ &+U(a_{1},..,a_{N})+\frac{c^{2}}{2}\sum_{n=1}^{N}\kappa_{n}=0,\end{split}$
@@
$\begin{split}L_{N}&=\frac{1}{2}\sum_{n=1}^{N}\dot{a}_{n}^{2}+\frac{4\pi G}{3}\sum_{n=1}^{N}\rho_{n}(a_{n})a_{n}^{2}-\\ &-\alpha^{2}\sum_{n=1}^{N}\left(\frac{1}{a_{n+1}-a_{n}}-\frac{1}{a_{n}-a_{n-1}}\right)^{2},\end{split}$
@@ Acknowledgments
The article was supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no. 075-02-2025-1789)